OpenStudy (anonymous):

I just have this last one @zilchero What is the slope of the line whose equation is –48 = 2x − 8y?

OpenStudy (anonymous):

y = mx + b get the equation into that form first

OpenStudy (anonymous):

-48=2x+8 ????????????????????

OpenStudy (anonymous):

Still there @zilchero?

OpenStudy (anonymous):

yes, but this thing has been glitching, so it's hard to type anything before i'm bumped back to the home page you cannot just erase variables. let's do some moving around -48 = 2x - 8y do you know how to move terms around? to get to 8y = 2x + 48 ?

OpenStudy (anonymous):

commutative property?

OpenStudy (anonymous):

Uh, sort of. I'll show you an example: 5 + x = 8 you need to get x alone on one side, so you need to get rid of the 5. but how? how do you get 5 to go away? subtract five. But since you do one thing to one side, you have to do the same thing to the other. 5 - 5 + x = 8 - 5 0 + x = 3 x = 3 that can be done with multiple variables as well -48 = 2x - 8y we want to isolate y this time, because y = mx + b let's get rid of 2x by subtracting 2x from both sides -48 - 2x = 2x - 2x - 8y -48 - 2x = 0 - 8y i think we can agree that 2 + 2 = 3 + 1 can also be written as 3 + 1 = 2 + 2. All i did was flip it. I'm going to do the same thing for -48 - 2x = 0 - 8y, since the concept is the same -8y = -2x -48 remember to keep the signs! anything that has a negative (or a minus sign) in front of it must be negative! i'll continue the problem in another post

OpenStudy (anonymous):

-8y = -2x - 48 we need to get y alone, as I said. So we need to get rid of the -8 in front of it. how? by dividing by eight. -8y is the same thing as -8 times y -8 divided by -8 is one, which means we got rid of the -8 so we need to divide by eight ON BOTH SIDES, because what you do to one side you must do to the other. -8y/-8 = -2/-8(x) - 48/-8

OpenStudy (anonymous):

can you simplify that for me?

OpenStudy (anonymous):

I think so hold on a sec

OpenStudy (anonymous):

-2/ -8(x) - 40???

OpenStudy (anonymous):

where did the y go?

OpenStudy (anonymous):

-2y

OpenStudy (anonymous):

Okay, wait a second

OpenStudy (anonymous):

 how about now?